# Two tales for a geometric Jensen–Shannon divergence

### v1

@@ 3.2 Closed-form formula for Gaussian distributions
Let us report the corresponding closed-form formula for Gaussian distributions:
@@
${\mathrm{JS}}^{G_{\alpha}}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{2},\Sigma_{2})})$ $=$ $(1-\alpha)\,D_{\mathrm{KL}}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{\alpha},\Sigma_{\alpha})})+\alpha\,D_{\mathrm{KL}}(p_{(\mu_{2},\Sigma_{2})},p_{(\mu_{\alpha},\Sigma_{\alpha})}),$
$=$ $(1-\alpha)\,B_{F}((\theta_{1}\theta_{2})_{\alpha},\theta_{1})+\alpha\,B_{F}((\theta_{1}\theta_{2})_{\alpha},\theta_{2}),$
$=$ $\frac{1}{2}\left(\mathrm{tr}\left(\Sigma_{\alpha}^{-1}((1-\alpha)\Sigma_{1}+\alpha\Sigma_{2})\right)+\log\left(\frac{|\Sigma_{\alpha}|}{|\Sigma_{1}|^{1-\alpha}|\Sigma_{2}|^{\alpha}}\right)\right$.
$=$ $\left.+(1-\alpha)(\mu_{\alpha}-\mu_{1})^{\top}\Sigma_{\alpha}^{-1}(\mu_{\alpha}-\mu_{1})+\alpha(\mu_{\alpha}-\mu_{2})^{\top}\Sigma_{\alpha}^{-1}(\mu_{\alpha}-\mu_{2})-d\right)$
${\mathrm{JS}}^{G_{\alpha}}_{*}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{2},\Sigma_{2})})$ $=$ $(1-\alpha)\,D_{\mathrm{KL}}(p_{(\mu_{\alpha},\Sigma_{\alpha})},p_{(\mu_{1},\Sigma_{1})})+\alpha D_{\mathrm{KL}}(p_{(\mu_{\alpha},\Sigma_{\alpha})},p_{(\mu_{2},\Sigma_{2})}),$
$=$ $(1-\alpha)\,B_{F}(\theta_{1},(\theta_{1}\theta_{2})_{\alpha})+\alpha\,B_{F}(\theta_{2},(\theta_{1}\theta_{2})_{\alpha}),$
$=$ $J_{F,\alpha}(\theta_{1},\theta_{2}),$
$=$ $\frac{1}{2}\left((1-\alpha)\mu_{1}^{\top}\Sigma_{1}^{-1}\mu_{1}+\alpha\mu_{2}^{\top}\Sigma_{2}^{-1}\mu_{2}-\mu_{\alpha}^{\top}\Sigma_{\alpha}^{-1}\mu_{\alpha}+\log\frac{|\Sigma_{1}|^{1-\alpha}|\Sigma_{2}|^{\alpha}}{|\Sigma_{\alpha}|}\right),,$
@@
$\Sigma_{\alpha}=\left((1-\alpha)\Sigma_{1}^{-1}+\alpha\Sigma_{2}^{-1}\right)^{-1},$ (43)
@@
$\mu_{\alpha}=\Sigma_{\alpha}\left((1-\alpha)\Sigma_{1}^{-1}\mu_{1}+\alpha\Sigma_{2}^{-1}\mu_{2}\right).$ (44)
4 Extended and normalized G-JSDs as regularizations of the JSD
The $M$-Jensen–Shannon divergence $\mathrm{JS}_{\mathrm{KL}}^{M}(p,q)$ can be interpreted as a regularization of the ordinary JSD:
@@ 4 Extended and normalized G-JSDs as regularizations of the JSD
$D_{\mathrm{JS}}^{M}(p_{1},p_{2})=D_{\mathrm{JS}}(p_{1},p_{2})+D_{\mathrm{KL}}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right).$ (45)
@@
$\mathrm{JS}_{\mathrm{KL}}^{M}(p_{1},p_{2})$ ${:=}$ $\frac{1}{2}\left(D_{\mathrm{KL}}(p_{1}:(p_{1}p_{2})_{M})+D_{\mathrm{KL}}(p_{2}:(p_{1}p_{2})_{M})\right),$
@@
$=$ $\frac{1}{2}\int\left(p_{1}\log\frac{p_{1}}{(p_{1}p_{2})_{A}}+p_{2}\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}+p_{2}\log\frac{p_{2}}{(p_{1}p_{2})_{A}}+p_{2}\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}\right)\,{\mathrm{d}\mu},$
@@
$=$ $D_{\mathrm{JS}}(p_{1},p_{2})+\int(p_{1}p_{2})_{A}\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}\,{\mathrm{d}\mu},$
$=$ $D_{\mathrm{JS}}(p_{1},p_{2})+D_{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M}).$
@@
Corollary 4 ($M$-JSD lower bound)
We have $D_{\mathrm{JS}}^{M}(p,q)\geq D_{\mathrm{JS}}(p,q)$.
@@
Since $D_{\mathrm{JS}}^{M}(p,q)=D_{\mathrm{JS}}(p,q)+D_{\mathrm{KL}}\left(\frac{p+q}{2},(pq)_{M}\right)$ and $D_{\mathrm{KL}}\geq 0$ by Gibbs’ inequality, we have $D_{\mathrm{JS}}^{M}(p,q)\geq D_{\mathrm{JS}}(p,q)$. $\square$
@@
Since the extended $M$-JSD is $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})=D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})+Z_{\alpha}-\log(Z_{\alpha})-1$, the extended $M$-JSD $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}$ can also be interpreted as another regularization of the Jensen–Shannon divergence when dealt with probability densities:
$D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})=D_{\mathrm{JS}}(p_{1},p_{2})+D_{\mathrm{KL}}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right)+Z_{\alpha}(p_{1},p_{2})-\log(Z_{\alpha}(p_{1},p_{2}))-1.$ (46)

### v3

@@ 3.2 Closed-form formula for Gaussian distributions
Let us report the corresponding closed-form formula for $d$-variate Gaussian distributions.
When $\alpha=\frac{1}{2}$, we proved that $\mathrm{JS}_{G}(p_{1},p_{2})=\frac{1}{4}\,J(p_{1},p_{2})-B(p_{1},p_{2})$ and $\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})=\frac{1}{4}\,J(p_{1},p_{2})+\exp(-B(p_{1},p_{2}))-1$ where $\mathrm{BC}(p_{1},p_{2})=\exp(-B(p_{1},p_{2}))$. Thus for the case of balanced geometric mixtures, we need to report the closed-form for the Jeffreys and Bhattacharyya distances:
$J(p_{\mu_{1},\Sigma_{1}},p_{\mu_{2},\Sigma_{2}})$ $=$ $\frac{1}{2}\,\left(\mathrm{tr}\left(\Sigma_{1}\Sigma_{2}^{-1}+\Sigma_{2}\Sigma_{1}^{-1}\right)+(\mu_{1}-\mu_{2})^{\top}(\Sigma_{1}^{-1}+\Sigma_{2}^{-1})(\mu_{1}-\mu_{2})-2d\right)\ ,$
$B(p_{\mu_{1},\Sigma_{1}},p_{\mu_{2},\Sigma_{2}})$ $=$ $\frac{1}{8}\,(\mu_{1}-\mu_{2})^{\top}\bar{\Sigma}^{-1}(\mu_{1}-\mu_{2})+\frac{1}{2}\log\left(\frac{\det\bar{\Sigma}}{\sqrt{\det\Sigma_{1}\,\det\Sigma_{2}}}\right),$
where $\bar{\Sigma}=\frac{1}{2}\,\left(\Sigma_{1}+\Sigma_{2}\right)$.
Otherwise, for arbitrary weighted geometric mixture $G_{\alpha}$, define $(\theta_{1}\theta_{2})_{\alpha}=\alpha\theta_{1}+(1-\alpha)\theta_{2}$, the weighted linear interpolation of the natural parameters $\theta_{1}$ and $\theta_{2}$.
@@
${\mathrm{JS}}_{G_{\alpha}}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{2},\Sigma_{2})})$ $=$ $\alpha\,{\mathrm{KL}}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{\alpha},\Sigma_{\alpha})})+(1-\alpha)\,{\mathrm{KL}}(p_{(\mu_{2},\Sigma_{2})},p_{(\mu_{\alpha},\Sigma_{\alpha})}),$
$=$ $\alpha\,B_{F}((\theta_{1}\theta_{2})_{\alpha},\theta_{1})+(1-\alpha)\,B_{F}((\theta_{1}\theta_{2})_{\alpha},\theta_{2}),$
$=$ $\frac{1}{2}\left(\mathrm{tr}\left(\Sigma_{\alpha}^{-1}(\alpha\Sigma_{1}+(1-\alpha)\Sigma_{2})\right)+\log\left(\frac{|\Sigma_{\alpha}|}{|\Sigma_{1}|^{\alpha}|\Sigma_{2}|^{1-\alpha}}\right)\right$.
$=$ $\left.+\alpha(\mu_{\alpha}-\mu_{1})^{\top}\Sigma_{\alpha}^{-1}(\mu_{\alpha}-\mu_{1})+(1-\alpha)(\mu_{\alpha}-\mu_{2})^{\top}\Sigma_{\alpha}^{-1}(\mu_{\alpha}-\mu_{2})-d\right)$
${\mathrm{JS}}_{G_{\alpha}}^{*}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{2},\Sigma_{2})})$ $=$ $(1-\alpha)\,{\mathrm{KL}}(p_{(\mu_{\alpha},\Sigma_{\alpha})},p_{(\mu_{1},\Sigma_{1})})+\alpha{\mathrm{KL}}(p_{(\mu_{\alpha},\Sigma_{\alpha})},p_{(\mu_{2},\Sigma_{2})}),$
$=$ $\alpha\,B_{F}(\theta_{1},(\theta_{1}\theta_{2})_{\alpha})+(1-\alpha)\,B_{F}(\theta_{2},(\theta_{1}\theta_{2})_{\alpha}),$
$=$ $J_{F,\alpha}(\theta_{1},\theta_{2})=:{B_{\alpha}}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{2},\Sigma_{2})}),$
$=$ $\frac{1}{2}\left(\alpha\mu_{1}^{\top}\Sigma_{1}^{-1}\mu_{1}+(1-\alpha)\mu_{2}^{\top}\Sigma_{2}^{-1}\mu_{2}-\mu_{\alpha}^{\top}\Sigma_{\alpha}^{-1}\mu_{\alpha}+\log\frac{|\Sigma_{1}|^{\alpha}|\Sigma_{2}|^{1-\alpha}}{|\Sigma_{\alpha}|}\right),$
@@
$\Sigma_{\alpha}=\left(\alpha\Sigma_{1}^{-1}+(1-\alpha)\Sigma_{2}^{-1}\right)^{-1},$ (45)
@@
$\mu_{\alpha}=\Sigma_{\alpha}\left(\alpha\Sigma_{1}^{-1}\mu_{1}+(1-\alpha)\Sigma_{2}^{-1}\mu_{2}\right).$ (46)
4 Extended and normalized G-JSDs as regularizations of the ordinary JSD
The $M$-Jensen–Shannon divergence $\mathrm{JS}_{M}(p,q)$ can be interpreted as a regularization of the ordinary JSD:
@@ 4 Extended and normalized G-JSDs as regularizations of the ordinary JSD
$\mathrm{JS}_{M}(p_{1},p_{2})=\mathrm{JS}(p_{1},p_{2})+{\mathrm{KL}}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right).$ (47)
@@
$\mathrm{JS}_{M}(p_{1},p_{2})$ ${:=}$ $\frac{1}{2}\left({\mathrm{KL}}(p_{1},(p_{1}p_{2})_{M})+{\mathrm{KL}}(p_{2},(p_{1}p_{2})_{M})\right),$
@@
$=$ $\frac{1}{2}\int\left(p_{1}\log\frac{p_{1}}{(p_{1}p_{2})_{A}}+p_{1}\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}+p_{2}\log\frac{p_{2}}{(p_{1}p_{2})_{A}}+p_{2}\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}\right)\,{\mathrm{d}\mu},$
@@
$=$ $\mathrm{JS}(p_{1},p_{2})+\int(p_{1}p_{2})_{A}\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}\,{\mathrm{d}\mu},$
$=$ $\mathrm{JS}(p_{1},p_{2})+{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M}).$
@@
Remark 13
One way to symmetrize the KLD is to consider two distinct symmetric means $M_{1}(a,b)=M_{1}(b,a)$ and $M_{2}(a,b)=M_{2}(b,a)$ and define
${\mathrm{KL}}_{M_{1},M_{2}}(p_{1},p_{2})={\mathrm{KL}}((p_{1}p_{2})_{M_{1}},(p_{1}p_{2})_{M_{2}})={\mathrm{KL}}_{M_{1},M_{2}}(p_{2},p_{1}).$
We notice that $\sqrt{{\mathrm{KL}}^{A,G}}$ is not a metric distance by reporting a triple of points $(p_{1},p_{2},p_{3})$ that fails the triangle inequality. Consider $p_{1}=(0.55,0.45)$, $p_{2}=(0.002,0.998)$, and $p_{3}=(0.045,0.955)$. We have $\sqrt{{\mathrm{KL}}_{M_{1},M_{2}}(p_{1},p_{2})}=0.5374165\ldots$, $\sqrt{{\mathrm{KL}}_{M_{1},M_{2}}(p_{1},p_{3})}=0.1759400\ldots$, and $\sqrt{{\mathrm{KL}}_{M_{1},M_{2}}(p_{3},p_{2})}=0.08485931\ldots$. The triangle inequality defect is
$\sqrt{{\mathrm{KL}}_{M_{1},M_{2}}(p_{1},p_{2})}-(\sqrt{{\mathrm{KL}}_{M_{1},M_{2}}(p_{1},p_{3})}+\sqrt{{\mathrm{KL}}_{M_{1},M_{2}}(p_{3},p_{2})})=0.2766171\ldots.$
We can also similarly symmetrize the extended KLD as follows:
${\mathrm{KL}}_{{\tilde{M}}_{1},{\tilde{M}}_{2}}^{+}(q_{1},q_{2})={\mathrm{KL}}^{+}((q_{1}q_{2})_{{\tilde{M}}_{1}},(q_{1}q_{2})_{{\tilde{M}}_{2}})={\mathrm{KL}}_{{\tilde{M}}_{1},{\tilde{M}}_{2}}(q_{2},q_{1}).$
In particular, when $M_{1}=A$ and $M_{2}=G$, we get the ${\mathrm{KL}}_{A,M}$ divergence:
${\mathrm{KL}}_{A,M}(p_{1},p_{2})=\frac{p_{1}+p_{2}}{2}\,\log\frac{p_{1}+p_{2}}{2\sqrt{p_{1}p_{2}}}+\log Z_{G}(p_{1},p_{2}),$
which is related to Taneja $T$-divergence [53]:
$T(p_{1},p_{2})=\int\frac{p_{1}+p_{2}}{2}\log\frac{p_{1}+p_{2}}{2\sqrt{p_{1}p_{2}}}.$ (48)
The $T$-divergence is a $f$-divergence [1, 13] obtained for the generator $f_{T}(u)=\frac{1+u}{2}\log\frac{1+u}{2\sqrt{u}}$.
Corollary 4 (JSD lower bound on $M$-JSD)
We have $\mathrm{JS}_{M}(p,q)\geq\mathrm{JS}(p,q)$.
@@
Since $\mathrm{JS}_{M}(p,q)=\mathrm{JS}(p,q)+{\mathrm{KL}}\left(\frac{p+q}{2},(pq)_{M}\right)$ and ${\mathrm{KL}}\geq 0$ by Gibbs’ inequality, we have $\mathrm{JS}_{M}(p,q)\geq\mathrm{JS}(p,q)$. $\square$
@@
Since the extended $M$-JSD is $\mathrm{JS}^{+}_{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})=\mathrm{JS}_{{M}_{\alpha},\beta}(p_{1},p_{2})+Z_{\alpha}-\log(Z_{\alpha})-1$, the extended $M$-JSD $\mathrm{JS}^{+}_{\tilde{M}_{\alpha},\beta}$ can also be interpreted as another regularization of the Jensen–Shannon divergence when dealt with probability densities:
$\mathrm{JS}^{+}_{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})=\mathrm{JS}(p_{1},p_{2})+{\mathrm{KL}}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right)+Z_{M_{\alpha}}(p_{1},p_{2})-\log(Z_{M_{\alpha}}(p_{1},p_{2}))-1.$ (49)
It is well-known that the JSD can be rewritten as a diversity index [32] using the concave entropy:
$\mathrm{JS}(p_{1},p_{2})=H\left(\frac{p_{1}+p_{2}}{2}\right)-\frac{H(p_{1})+H(p_{2})}{2}.$ (50)
We generalize this decomposition as a difference of a cross-entropy term minus entropies as follows:
Proposition 10 (M-JSD cross-entropy decomposition)
We have
$\mathrm{JS}_{M}(p_{1},p_{2})=H^{\times}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M})-\frac{H(p_{1})+H(p_{2})}{2}.$
Proof:
We have from Proposition 9:
$\mathrm{JS}_{M}(p_{1},p_{2})=\mathrm{JS}(p_{1},p_{2})+{\mathrm{KL}}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right).$
Since ${\mathrm{KL}}(p_{1},p_{2})=H^{\times}(p_{1},p_{2})-H(p_{1})$ where $H^{\times}(p_{1},p_{2})=-\int p_{1}\log p_{2}\,{\mathrm{d}\mu}$ is the cross-entropy and $H(p)=-\int p\log p\,{\mathrm{d}\mu}=H^{\times}(p,p)$ is the entropy. Pluggin Eq. 50 in Eq. 47, we get
$\mathrm{JS}_{M}(p_{1},p_{2})$ $=$ $H\left(\frac{p_{1}+p_{2}}{2}\right)-\frac{H(p_{1})+H(p_{2})}{2}+H^{\times}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right)-H\left(\frac{p_{1}+p_{2}}{2}\right),$
$=$ $H^{\times}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right)-\frac{H(p_{1})+H(p_{2})}{2}.$
Note that when $M=A$, the arithmetic mean, we have $H^{\times}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right)=H\left(\frac{p_{1}+p_{2}}{2}\right)$ and we recover the fact that $\mathrm{JS}_{M}(p_{1},p_{2})=\mathrm{JS}(p_{1},p_{2})$. $\square$
5 Estimation and approximation of the extended and normalized M-JSDs
Let us recall the two definitions of the extended M-JSD and the normalized M-JSD (for the case of $\alpha=\beta=\frac{1}{2}$) between two normalized densities $p_{1}$ and $p_{2}$:
$\mathrm{JS}_{M}(p_{1},p_{2})$ $=$ $\frac{1}{2}\left({\mathrm{KL}}\left(p_{1},(p_{1}p_{2})_{M}\right)+{\mathrm{KL}}\left(p_{2},(p_{1}p_{2})_{M}\right)\right),$
[…]
where $(p_{1}p_{2})_{M}(x)=\frac{M(p_{1}(x),p_{2}(x))}{Z_{M}(p_{1},p_{2})}$ (with $Z_{M}(p_{1},p_{2})=\int M(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}(x)$) and $(p_{1}p_{2})_{{\tilde{M}}}(x)=M(p_{1}(x),p_{2}(x))$.
In practice, one needs to estimate the extended and normalized G-JSDs when they do not admit closed-form formula.
5.1 Monte Carlo estimators
To estimate $\mathrm{JS}_{M}(p_{1},p_{2})$, we can use Monte Carlo samplings to estimate both KLD integrals and mixture normalizers $Z_{M}$; For example, the normalizer $Z_{M}(p_{1},p_{2})$ is estimated by
$\hat{Z}_{M}(p_{1},p_{2})=\frac{1}{s}\sum_{i=1}^{s}\frac{1}{r(x_{i})}\,M(p_{1}(x_{i}),p_{2}(x_{i})),$
where $r(x)$ is the proposal distribution which can be chosen according to the mean $M$ and the types of probability distributions $p_{1}$ and $p_{2}$, and $x_{1},\ldots,x_{s}$ are $s$ identically and independently samples (iid.) from $r(x)$. However, since $(p_{1}p_{2})_{M}(x)$ is now estimated as $(p_{1}p_{2})_{\hat{M}(x)}$, it is not anymore a normalized $M$-mixture, and thus we consider estimating
[…]
to ensure the non-negativity of the divergence $\mathrm{JSD}_{\hat{M}}$.
Let us consider the estimation of the term
${\mathrm{KL}}^{+}\left(p_{1},(p_{1}p_{2})_{{\tilde{M}}}\right)=\int\left(p_{1}\log\frac{p_{1}}{M(p_{1},p_{2})}+M(p_{1},p_{2})-p_{1}\right)\,{\mathrm{d}\mu}.$
[…]
$\widehat{{\mathrm{KL}}}^{+}\left(p_{1},(p_{1}p_{2})_{{\tilde{M}}}\right)=\frac{1}{s}\sum_{i=1}^{s}\left(\log\frac{p_{1}(x_{i})}{M(p_{1}(x_{i}),p_{2}(x_{i}))}+\frac{1}{p_{1}(x_{i})}\,M(p_{1}(x_{i}),p_{2}(x_{i}))-1\right).$
Monte Carlo (MC) stochastic integration [47] is a well-studied topic in Statistics with many results on consistency and variance of MC estimators.
Note that even if we have a generic formula for the G-JSD between two densities of an exponential family given by Corollary 2, the cumulant function $F(\theta)$ may not be in closed form [11, 24]. This is the case when the sufficient statistic vector of the exponential family is $t(x)=(x,x^{2},\ldots,x^{m})$ (for $m\geq 5$) yielding the polynomial exponential family (also called exponential-polynomial family [24]).
5.2 Approximations via $\gamma$-divergences
One way to circumvent the lack of computational tractable density normalizers is to consider the family of $\gamma$-divergences [19] instead of the KLD:
${\tilde{D}}_{\gamma}(q_{1},q_{2})=\frac{1}{\gamma(1+\gamma)}\log I_{\gamma}(q_{1},q_{2})-\frac{1}{\gamma}\log I_{\gamma}(q_{1},q_{2})+\frac{1}{1+\gamma}\log I_{\gamma}(q_{1},q_{2}),\quad\gamma>0,$
where
$I_{\gamma}(q_{1},q_{2})=\int q_{1}(x)\,q_{2}^{\gamma}(x)\,{\mathrm{d}\mu}(x).$
The $\gamma$-divergences are projective divergences, i.e., they enjoy the property that
${\tilde{D}}_{\gamma}(\lambda_{1}q_{1},\lambda_{2}q_{2})={\tilde{D}}_{\gamma}(q_{1},q_{2}),\quad\forall\lambda_{1}>0,\lambda_{2}>0.$
Furthermore, we have $\lim_{\gamma\rightarrow 0}{\tilde{D}}_{\gamma}(p_{1},p_{2})={\mathrm{KL}}(p_{1},p_{2})$. (Note that KLD is not projective.)
Let us define the projective $M$-JSD:
[…]
We have for $\gamma=\epsilon$ small enough (e.g., $\epsilon\leq 10^{-3}$), $\mathrm{JS}_{M}(p_{1},p_{2})\approx\mathrm{JS}_{{\tilde{M}},\gamma}(p_{1},p_{2})$ since
[…]
In particular, for exponential family densities $p_{\theta_{1}}(x)=\frac{q_{\theta_{1}}(x)}{Z(\theta_{1})}$ and $p_{\theta_{2}}(x)=\frac{q_{\theta_{2}}(x)}{Z(\theta_{2})}$, we have
$I_{\gamma}(p_{\theta_{1}},p_{\theta_{2}})=\exp\left(F(\theta_{1}+\gamma\theta_{2})-F(\theta_{1})-\gamma F(\theta_{2})\right),$
provided that $\theta_{1}+\gamma\theta_{2}$ belongs to the natural parameter space (otherwise, the integral $I_{\gamma}$ diverges).
Even when $F(\theta)$ is not known in closed form, we may estimate the $\gamma$-divergence by estimating the $I_{\gamma}$ integrals as follows:
$\hat{I}_{\gamma}(q_{\theta_{1}},q_{\theta_{2}})\approx\frac{1}{s}\sum_{i=1}^{s}q_{2}(x_{i}),$
where $x_{1},\ldots,x_{s}$ are iid. sampled from $p_{1}(x)$. For example, we may use Monte Carlo importance sampling methods [30] or exponential family Langevin dynamics [5] to sample densities of exponential family densities with computationally intractable normalizers (e.g., polynomial exponential families).